OpenStudy (anonymous):

You invest an initial $2,000 in an account that has an annual interest rate of 6% compunded daily. How much money will you have in the account after 15 years?

jimthompson5910 (jim_thompson5910):

Use the same formula as last time, but now P = 2000, r = 0.06, n = 365 (assuming it's not a leap year) and t = 15

jimthompson5910 (jim_thompson5910):

So you should have A = 2000(1+0.06/365)^(365*15)

OpenStudy (anonymous):

So to start: A= 2000(0.0029041095890411)^5475

jimthompson5910 (jim_thompson5910):

no, you divide 0.06 by 365 first before you add

jimthompson5910 (jim_thompson5910):

so it should be 1+0.06/365 1 + 0.00016438356164 1.00016438356164 So the overall expression should be A = 2000(1.00016438356164)^5475

jimthompson5910 (jim_thompson5910):

let me know what you get for the answer

OpenStudy (anonymous):

2.45942119*2000= $4,919

jimthompson5910 (jim_thompson5910):

very good, you got it

OpenStudy (anonymous):

Yay! =) Thank you so much for all your help! There are only three more problems I need some assistance on

jimthompson5910 (jim_thompson5910):

alright post them over to the left and I'll help you out